OpenStudy (anonymous):

What is the vertex of the graph of y=-|3x-12|+5?

OpenStudy (anonymous):

The x co-ordinate of the vertex of an absolute value graph will be the value of x which makes the expression within the absolute value signs equal to zero. The y value is what you get when you ignore the absolute value part of the equation. In this case, the x-value of the vertex is where 3x-12=0, so x=4. The y value is where y=5. Do you follow?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

If it makes it more clear to you, the vertex will be where the graphs of: \[y=-(3x-12)+5\] and \[y=3x-12+5\] cross. So if we equate them we get: \[3x-7=-3x+17\] \[6x=24\] \[x=4\] Substituting this into any of the equations will give the corresponding y-value: \[y=-3((4)-12)+5=5\]